A vertical rod of length l is moved with constant velocity v towards east. The vertical component of earth magnetic field is B and angle of dip is theta. What is the induced emf in the rod?

Grade:12th pass

4 Answers

Ritik
6 years ago
Induced EMF is given by E= Blv ......1 Vertical component of B is B sin theta From 1 ,we getE= (B sin theta)× L× v cos theta E= B/v cot theta
1234567890
5 years ago
This petition is made in regards to Hip Hop international Artist Daniel Hernandez stage name (Tekashi 69) His Family , Fans , Friends & Supporters hope that the judge can grant this young man bail so he can get back with his family & continue to make great music for his fans while he fights this legal battle. Let’s show that this Rainbow hair rapper is a non threat to any community but is loved for all that he has done & the positive impact he has on others. We thank everyone for your love & support 
ankit singh
3 years ago
A vertical conducting rod of length l is moved at a constant velocity v eastward. The vertical component of the Earth's magnetic field is B and the angle of dip is theta. ... The rod is vertical and moves east ward, therefore it 'cuts' horizontal component Bh of the magnetic field.
Naman Agrawal
2 years ago

From the above figure, are conclude that magnetic force which is responsible for induction of emf is horizontal component of magnetic field only.

 
Formula Δe=(Vˉ×Bˉ)l=VBlsin90=VBl
 
Now tanθ=BBHBH=Bcotθ
 
Now, emf induced = 
e=V(BHl)
 
e=VBlcotθ
 
e=Blvcotθ